# revealing an in-context neural scaling law

### v1

@@ 4.2.1 Brownian motion
4.2.2 Geometric Brownian motion
@@
4.2.3 Noisy logistic map
The logistic map system becomes stochastic when one introduces small Gaussian perturbations of variance $\sigma^{2}$ at each step to model noisy measurements: $x_{t+1}=f(x_{t}+\epsilon)$ where $\epsilon\stackrel{{\scriptstyle\text{i.i.d}}}{{\sim}}\mathcal{N}(0,\sigma^{2})$. This implies that the conditional distribution has non-zero uncertainty and is no longer delta-like [^2] The approximation in Equation 7 assumes a small perturbation compared to the second derivative, that is $\sigma^{2}\ll 1/f^{\prime\prime}(x)$.:
The first derivative of $f$ measures how sensitive the local dynamics are to external perturbations. This intuitively explains why the standard deviation of the conditional distribution should be proportional to $f^{\prime}$.
Figure 12: Stochastic logistic map in-context loss curves.
We observe a fast decay of the in-context loss function with respect to the length of the observed time series in Figure 12. This is in agreement with previous numerical experiments and suggests that the LLM’s prediction ability is robust against noise perturbation. Moreover, the correct model for the standard deviation of the time series is encoded in the LLM PDF, as shown in Figure 13, and can be extracted with the Hierarchy-PDF algorithm.
4.2.4 Markov chains with discrete states
$P_{ij}=P(X_{t+1}=j|X_{t}=i),\quad 1\leq i,j\leq n.$
Using the testing procedures elaborated in Section 3.2, we generate ten Markov chains, each from a distinct and randomly generated transition matrix of size $n=4$.
Figure 14: Markov chain in-context loss curves decay rapidly with respect to the input time series length.
The corresponding loss curves between the LLM predictions and the ground truth are displayed in Figure 14. Unlike the previous dynamical systems investigated earlier, which are all motivated by models pertaining to physical phenomena, the discrete Markov chains are commonly used to describe the evolution of symbolic systems, such as language (Almutiri & Nadeem, 2022) and decision-making process (Feinberg & Shwartz, 2012). In our experiments, the transition rules are synthesized completely at random. Yet, the LLM formulates remarkably accurate statistical predictions as more data are observed in-context. These conclusions hold for larger transition matrices ($n>4$) and bigger LLMs, such as LLaMA-70b (see Section A.5).
5 Discussion and conclusion
We discuss the limitations, interpretations, and implications of the experimental findings presented in this work.
In-context neural scaling law. Neural scaling laws (Kaplan et al., 2020) are power laws that characterize how the loss of trained neural networks vary with respect to parameters of the model, such as model size, dataset size, and computational resources. To the best of our knowledge, neural scaling laws have, so far, only been observed in the training procedure, which explicitly updates the weights of neural networks using optimization algorithms. The loss curves observed in the different numerical experiments (see Figure 15 in Section A.1) reveal an additional in-context scaling law for LLMs zero-shot learning of dynamical systems.
Early plateauing and stationary distributions. While most of our experimental datasets give rise to power scaling laws, two loss curves —the Brownian motion and geometric Brownian motion— “plateau early” at a context length of approximately $10^{2}$ (see Figure 15). One potential explanation for this phenomena, elaborated in Section A.2, is that the Brownian and geometric Brownian motions “wander out of distribution” at large time $t>0$, while all the other time series associated with the dynamical systems studied in this paper converge to stationary distributions.
Temperature. The temperature hyperparameter $T$ affects the outcome of the LLM softmax layer and, hence, the reconstructed PDF obtained through the Hierarchy-PDF algorithm. The next digits prediction are extracted at the parameter $T=1$, which is the default temperature used during training (Touvron et al., 2023). However, we found that using a temperature of $T=0.7$ might be more suitable for extracting predicted specific statistical parameters such as the standard deviation from the PDF (see Figures 11 and 13, and the discussion in Section A.3).
Data leakage. The possibility that LLMs’ accurate predictions of the next time step value are merely due to memorization seems extremely unlikely. While we cannot rule out the existence of numerical sequences in the training corpus (especially popular ones like the Fibonacci sequence or the digits of $\pi$), our experiments rely on custom time series with arbitrary parameters. In addition, these sequences have been normalized for proper tokenization. A sequence of even one thousand numerical values, encoded with three digits, may cover $10^{3000}$ particular instances, which is well beyond the $\sim 10^{12}$ tokens of the training corpus (Touvron et al., 2023). Therefore, data leakage in the context of numerical time series appears implausibly unlikely.
In conclusion, we showed that, with increasing cardinality in tokenization, and given sufficient context, LLMs can accurately extrapolate not just deterministic trajectories (non-linear pendulum), but also chaotic and stochastic sequences governed by specific transition rules. In the latter case, the extrapolation accuracy is measured in a statistical sense. This suggests that large language models can in-context learn dynamical systems’ time series and predict future states in a manner that maintains fidelity with the underlying principles governing the system’s evolution. Moreover, this behavior is observed without the use of any fine-tuning or “prompting engineering”.
The implications are substantial. Our results suggest that a transformer, trained primarily on textual data, can extract governing principles of numerical sequences observed in-context. This further supports the notion that LLMs may somehow encode in their architecture, a rich and detailed representation of the world. Prior investigations of the “world model” hypothesis (Gurnee & Tegmark, 2023; Li et al., 2023; Patel & Pavlick, 2021) primarily focus on internal representations of static semantic concepts. In contrast, our study indicates that an LLM’s world model also encompasses the mathematical principles underlying the dynamics of both physical systems, like the logistic map, and symbolic systems, such as Markov chains.
An exciting future direction is to study the generalization of the observed in-context neural scaling laws for other LLMs, such as GPT4 (OpenAI, 2023), and the newly proposed state space models (Gu & Dao, 2023). Furthermore, the in-context neural scaling law might hint at a learning algorithm that LLMs implicitly implement during inference (e.g., a numerical time-stepping scheme), and characterizing such an algorithm is an open question of broad interest.
Software and Data
Broader Impact
While there are several potential societal consequences of large language models, the main goal of this work is to advance the field of machine learning.
Acknowledgements
This work was supported by the SciAI Center, and funded by the Office of Naval Research (ONR), under Grant Numbers N00014-23-1-2729 and N00014-23-1-2716.
Appendix A Appendix
A.1 In-context neural scaling law
@@
$L(N)$ $=\left(\frac{N}{N_{\text{c}}}\right)^{\alpha_{N}},$ (8)
$L(D)$ $=\left(\frac{D}{D_{\text{c}}}\right)^{\alpha_{D}},$ (9)
$L(C)$ $=\left(\frac{C}{C_{\text{c}}}\right)^{\alpha_{C}},$ (10)
where $L$ represents the loss or performance metric of the model. The characteristic factors ($N_{\text{c}}$, $D_{\text{c}}$, and $C_{\text{c}}$), and power coefficient ($\alpha$) are extracted empirically from training curves. The fitted quantities depend on the distribution of data, the model architecture and the type of optimizer used for training. Such power-law relations appear in log-log plots as straight lines, whose slopes correspond to the parameter $\alpha$.
Our loss curves from learning dynamical systems (see Figure 2) reveal an additional neural scaling law that applies to in-context learning:
$L(D_{\text{in}})=\left(\frac{D_{\text{in}}}{D_{\text{c}}}\right)^{\alpha},$ (11)
@@
To the best of our knowledge, neural scaling laws have so far only been observed in the training procedure, which updates the weights of neural networks using an explicit algorithm, such as stochastic gradient descent and Adam (Kingma & Ba, 2017).
A.2 Invariant measure and the early plateauing of in-context loss
While most datasets are well-described by the power laws, two loss curves — the Brownian motion and geometric Brownian motion — plateau early at a context length of about $10^{2}$, as shown in Figure 15. We attribute this early plateauing to the fact that the Brownian and geometric Brownian motions “wander out of distribution” at large $t$, while all other dynamical systems studied in this paper converge to stable distributions (i.e., the invariant measure).
A Markovian system (stochastic or deterministic) governed by a transition rule $P(x_{t+1}|x_{t})$ is said to have an invariant measure $\pi(x)$ if
[…]
@@ A.2 Invariant measure and the early plateauing of in-context loss
which is analogous to the continuous case described by Equation 12. By definition, any non-negative right eigenvector of $p$ with eigenvalue $\lambda=1$ is a stationary distribution of $p$. (Sethna, 2021) showed that a valid transition matrix has at least one stationary distribution.
On the other hand, neither the Brownian nor the geometric Brownian motion has invariant distributions [^3] For stochastic systems, the invariant measure is sometimes referred to as the stationary distribution. on unbounded domains (e.g., when $\mathcal{X}=\mathbb{R}$). This can be seen from the marginalized distribution $P(x_{t})$ at time $t$. For the Brownian motion defined in Equation 4, the marginalized distribution of $x_{t}$ at time $t$ is a normal distribution:
@@
Both Equations 14 and 15 are time-dependent and do not converge to a stationary distribution in the limit $t\rightarrow\infty$.
For the Brownian and geometric Brownian motions, the LLM might decide to only consider the most recent segment of time steps, and ignore the earlier data, which are in some sense “out of distribution”. This could explain the early plateauing of loss curves. Indeed, the classical neural scaling laws can be improved or broken if the scheduling of the training data shifts in distribution, as shown in (Sorscher et al., 2023). Different from (Sorscher et al., 2023; Lu et al., 2023), which alter the scheduling of data to achieve better learning curves that decrease faster with the size of training data, our experiments consider time series with pre-determined transition laws. We cannot tamper with the scheduling of our data to make it stationary without altering the underlying transition rules.
A.3 Temperature and variance
The temperature $T$ is a hyper-parameter that controls the variance of the softmax output layer. Although most LLMs are trained at $T=1$, it is common practice to tune the temperature in the interval $T\in[0.8,1.2]$ during inference. Then, one can opt for increased diversity (high $T$), or better coherence (low $T$) in the generated output. The temperature hyper-parameter affects the uncertainty, or variance, in the Hierarchy-PDF extracted from the LLM. Figures 17 and 17 show how different temperatures change the shape of the Hierarchy-PDF. In both cases, higher temperature leads to higher variance in the PDF.
@@ A.3 Temperature and variance
While the loss curves in our paper are calculated at $T=1$, the predicted $\sigma$ shown in Figures 11 and 13 are extracted at $T=0.7$. We performed a grid search on the temperature ranging from $T=0.3$ to $T=3$ (see Figures 20 and 20), and observed that $T=0.7$ consistently results in better prediction quality of the variance.
@@
A.4 Non-linear pendulum
A non-linear pendulum is a classical physical system that extends the simple pendulum model by considering the effects of large angular displacements. Unlike the linear approximation, which assumes small oscillations and uses a simple harmonic oscillator equation, the non-linear pendulum model incorporates the full trigonometric term, which leads to complex non-linear phenomena such as amplitude-dependent frequency:
$\ddot{\theta}=-\lambda\dot{\theta}-\omega^{2}\sin(\theta).$ (16)
In the simulation shown in Figure 1, we use the initial condition $(\theta,\dot{\theta})=(0,1.05)$, the natural frequency $\omega=2$, and damping coefficient $\lambda=0.1$. The non-zero damping results in both amplitude and frequency drifting over time.
A.5 Markov chains with larger numbers of states and bigger LLM
A.6 Additional loss functions
Although commonly used as the training loss for a variety of machine learning systems, this loss function may suffer from numerical instabilities as the learned transition function $\tilde{P}$ are often close to zero, as shown in Figure 5 and Figure 8, where the probability density is concentrated in small regions of the support.
$=-\frac{1}{2}\ln(\tilde{p}(x_{\mathrm{true}}))+\text{constant}.$
As a result, the Bhattacharyya distance is reduced to an affine-transformed negative log-likelihood assigned to data by the LLM. Such local sensitivity on $\tilde{p}(x_{\mathrm{true}})$ explains the wild fluctuations seen in the Bhattacharyya loss in Section 4.1.
The kurtosis is equal to $3$ for a Gaussian distribution, and $\frac{9}{5}$ for bounded uniform distributions. Figure 22 shows the kurtosis of Brownian motion transition rules learned by LLM, which indeed converges to $3$ with increasing context length.
A.7 Hierarchy PDF

### v2

@@ A.3.1 Brownian motion
A.3.2 Geometric Brownian motion
@@
A.3.3 Markov chains with LLaMA-70b
A.4 Additional Experiments: deterministic time series
A.4.1 Logistic map
$x_{t+1}=f(x_{t})=rx_{t}(1-x_{t}),\quad x_{0}\in(0,1),$ (11)
which may also be written using conditional distributions to reflect the deterministic nature of the system as $X_{t+1}|\{X_{t}=x_{t}\}\sim\delta_{f(x_{t})}$, where $\delta$ denotes the Dirac delta distribution. This conditional distribution is the ground truth transition function displayed in red in Figure 13. The parameter $r\in[1,4)$ controls the behavior of the system and is set to $r=3.9$. At this value, the dynamics are naturally confined within the interval $(0,1)$, and the system has no stable fixed points. Due to the chaotic nature of the system, two initial nearby trajectories diverge exponentially in time. This property allows us to sample multiple uncorrelated trajectories by using different initial conditions $x_{0}$, sampled uniformly in $(0,1)$.
A.4.2 Lorenz system
$\dot{x}(t)=\sigma(y-x),\quad\dot{y}(t)=x(\rho-z)-y,\quad\dot{z}(t)=xy-\beta z,$
A.5 Continuous State Space Visualization
One may naively remark upon the possibility that the in-context learning task for the Lorenz system and the logistic map could be rendered trivial if $x_{t+1}$ always falls close to $x_{t}$, in which case the LLM only needs to learn a static noisy distribution in order to decrease the loss. This is not the case with our experiments. In this section, we demonstrate the non-triviality of the learning tasks in Figures 17 and 17. In both cases, it is clear that the LLM has successfully learned to actively predict the expected mean position of the next state, and, in the logistic map example, the variance of the next state distribution as well. We note that the Lorenz system is simulated deterministically, hence the true next-state distribution is represented as a delta-function.
Figure 16: 4 consecutive states in a noisy logistic map. Ground truth is shown in red and LLaMA predictions in blue.
Figure 17: 4 consecutive states in a Lorenz system trajectory. Ground truth is shown in red and LLaMA predictions in blue.
A.6 In-context neural scaling law
@@
$L(N)=\left(\frac{N}{N_{\text{c}}}\right)^{\alpha_{N}},\quad L(D)=\left(\frac{D}{D_{\text{c}}}\right)^{\alpha_{D}},\quad L(C)=\left(\frac{D}{C_{\text{c}}}\right)^{\alpha_{C}},$
where $L$ represents the loss or performance metric of the model. The characteristic factors ($N_{\text{c}}$, $D_{\text{c}}$, and $C_{\text{c}}$), and power coefficient ($\alpha$) are extracted empirically from training curves. The fitted quantities depend on the distribution of data, the model architecture, and the type of optimizer used for training. Such power-law relations appear in log-log plots as straight lines, whose slopes correspond to the parameter $\alpha$. Our loss curves from learning dynamical systems (see Figure 1) reveal an additional neural scaling law that applies to in-context learning:
$L(D_{\text{in}})=\left(\frac{D_{\text{in}}}{D_{\text{c}}}\right)^{\alpha},$
@@
A.7 Baselines for noisy logistic map and Markov chains
In this section, we compare LLM’s predictions against baseline models of known architectures, in order to understand the difficulty of the in-context learning task and make better sense of the Bhattacharyya loss. Specifically, we consider the following baseline models: unigram and bi-gram models for discrete Markov chains, and linear and non-linear autoregressive models with 1-step memory (AR1) for noisy logistic maps. The bi-gram model for the Markov chain has an unfair advantage since it is designed to model Markovian processes where the probability distribution of a token depends only on the previous token, ie., inferring the values of the transition matrix. The unigram model, on the other hand, models all tokens as drawn i.i.d. from the same distribution.
Figure 19: LLM in-context loss curves against the baseline model loss curves. The coefficient $\alpha$ denotes the fitted scaling exponent.
The neural network AR1 model takes a state $x_{t-1}$ as input, and outputs prediction for next state $x_{t}$ as a Gaussian distribution parameterized by mean and variance: $f_{\theta}:x_{t}\rightarrow\mathcal{N}(\mu_{\theta}(x_{t}),\sigma_{\theta}(x_{t}))$. As such, it also has the unfair advantage of hard-coded Gaussianiety. LLaMA, on the other hand, must infer the correct distribution family from data. Despite this intrinsic disadvantage, LLaMA still outperforms the neural network AR1 model in the large context limit. The NN used in the non-linear AR1 models features three fully connected hidden layers of widths 64, 32, and 16. We found that simpler neural networks are easily trapped in local minima, leading to unstable performance. The loss curves in Figure 19 are obtained by training an independent copy of this neural network to convergence at each context length, and predict the next state distribution using the trained NN. The training loss is defined as the negative log-likelihood of the observation data:
$\mathcal{L}(\textup{data},\theta)$ $=-\sum_{{x_{t-1},x_{t}}\in\textup{data}}\log P(x_{t};\mu_{\theta}(x_{t-1}),\sigma_{\theta}(x_{t-1}))$
$=\sum_{{x_{t-1},x_{t}}\in\textup{data}}\log\sigma_{\theta}(x_{t-1})-\frac{1}{2}\left(\frac{x_{t}-\mu_{\theta}(x_{t-1})}{\sigma_{\theta}(x_{t-1})}\right)^{2}.$
Furthermore, the ensemble of NNs allows us to visualize the learned transition functions $P(x_{t+1}|x_{t})$ at each context length. In Figure 20, we show how the transition rules learned by the NNs gradually converge to the ground truth as context length increases. Since at large context length, the LLMs achieve similar loss as the NN-based AR1 model, it is reasonable to expect the LLM to have learned a transition function of similar accuracy as shown in the 5th plot in Figure 20. However, it is difficult to visualize the full transition rules $P(x_{t+1}|x_{t})$, for $x_{t}\in[0,1]$, as learned by an LLM, because doing so would require appending an array of $x_{t}$s at the end of a training sequence, which would render the training sequence incorrect.
Figure 20: Noisy logistic map transition rules, $P(x_{t}|x_{t}-1)$, learned by a neural network-based AR1 model against the ground truth transition rule.
A.8 Invariant measure and the early plateauing of in-context loss
While most datasets are well-described by the power laws, two loss curves — the Brownian motion and geometric Brownian motion — plateau early at a context length of about $10^{2}$, as shown in Figure 18. We attribute this early plateauing to the fact that the Brownian and geometric Brownian motions “wander out of distribution” at large time $t$, while all other dynamical systems studied in this paper converge to stable distributions (i.e., the invariant measure). A Markovian system (stochastic or deterministic) governed by a transition rule $P(x_{t+1}|x_{t})$ is said to have an invariant measure $\pi$ if
$\pi(x_{t+1})=\int_{\mathcal{X}}\pi(x_{t})P(x_{t+1}|x_{t})\,\textup{d}x_{t},\quad x_{t+1}\in\mathcal{X}.$ (12)
@@ A.8 Invariant measure and the early plateauing of in-context loss
which is analogous to the continuous case described by Equation 12. By definition, any non-negative right eigenvector of $p$ with eigenvalue $\lambda=1$ is a stationary distribution of $p$. (Sethna, 2021) showed that a valid transition matrix has at least one stationary distribution. On the other hand, neither the Brownian nor the geometric Brownian motion has invariant distributions [^3] For stochastic systems, the invariant measure is sometimes referred to as the stationary distribution. on unbounded domains (e.g., when $\mathcal{X}=\mathbb{R}$). This can be seen from the marginalized distribution $P(x_{t})$ at time $t$. For the Brownian motion defined in Equation 8, the marginalized distribution of $x_{t}$ at time $t$ is a normal distribution:
@@
Both Equations 14 and 15 are time-dependent and do not converge to a stationary distribution in the limit $t\rightarrow\infty$. For the Brownian and geometric Brownian motions, the LLM might decide to only consider the most recent segment of time steps, and ignore the earlier data, which are in some sense “out of distribution”. This could explain the early plateauing of loss curves. Indeed, the classical neural scaling laws can be improved or broken if the scheduling of the training data shifts in distribution, as shown in (Sorscher et al., 2023). Different from (Sorscher et al., 2023; Lu et al., 2022), which alter the scheduling of data to achieve better learning curves that decrease faster with the size of training data, our experiments consider time series with pre-determined transition laws. We cannot tamper with the scheduling of our data to make it stationary without altering the underlying transition rules.
A.9 Temperature and variance
The temperature $T$ is a hyper-parameter that controls the variance of the softmax output layer. Although most LLMs are trained at $T=1$, it is common practice to tune the temperature in the interval $T\in[0.8,1.2]$ during inference. Then, one can opt for increased diversity (high $T$), or better coherence (low $T$) in the generated output. The temperature hyper-parameter affects the uncertainty, or variance, in the Hierarchy-PDF extracted from the LLM. Figures 22 and 22 show how different temperatures change the shape of the Hierarchy-PDF. In both cases, higher temperature leads to higher variance in the PDF.
@@ A.9 Temperature and variance
While the loss curves in our paper are calculated at $T=1$, the predicted $\sigma$ shown in Figures 11 and 6 are extracted at $T=0.7$. We performed a grid search on the temperature ranging from $T=0.3$ to $T=3$ (see Figures 25 and 25), and observed that $T=0.7$ consistently results in better prediction quality of the variance.
@@
A.10 Hierarchy PDF